Find the magnitude of the charge if it has a Voltage of 13 V with a distance of 25 mm


1
Expert's answer
2022-03-20T18:50:42-0400

The voltage produced by a point charge

"V=k\frac{q}{r}"

Hence, the magnitude of charge

"q=\frac{Vr}{k}=\frac{13\times 0.025}{9\times 10^9}=36\times 10^{-12}\:\rm C=36\: pC"